OpenStudy (hockeychick23):

Pre-calc help please? A certain radioactive isotope is a by-product of some nuclear reactors. Due to an explosion, a nuclear reactor experiences a massive leak of this radioactive isotope. Fortunately, the isotope has a very short half-life of 10 days. After 2 days, about what percent of the original amount of the isotope released by the explosion remains?

OpenStudy (hockeychick23):

@Directrix @ILovePuppiesLol could you help me please?

OpenStudy (unklerhaukus):

The number of atoms \(N\) as a function of time past \(t\), is equal to the initial number \(N_0\) times a half to the power of the time past, divided by the half-life t1/2. \[\large N(t)=N_0(\tfrac12)^{\frac t{t_{1/2}}}\] The ratio \(N(t)/N_0\) is the ratio of the remaining isotope after a time \(t\). So. substitute \(t=2\), and \(t_{1/2}=10\), into: \[\large\frac{N(t)}{N_0}=(\tfrac12)^{\frac t{t_{1/2}}}\]

OpenStudy (unklerhaukus):

@hockeychick23